OpenStudy (anonymous):

Solve the system by substitution. -4.5x - 2y = -12.5 3.25x - y = -0.75

OpenStudy (aaronq):

oi wrong forum (1) -4.5x - 2y = -12.5 isolate y= (12.5 -4.5x)/2 substitute into (2) 3.25x - y = -0.75 3.25x-((12.5-4.5x)/2) = -0.75 6.5x-12.5+4.5x=-1.5 11x=11 x=1 substitute into (1) y= (12.5 -4.5x)/2 y=(12.5-4.5(1))/2 y=4

OpenStudy (anonymous):

opps didn't know i put it here but thanks for the answer :)

OpenStudy (aaronq):

no prob dude !